Circular No. 8754

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006gk–2006gz Ten additional type-Ia supernovae (2006gk–2006gp, 2006gu–2006gx)

and one probable type-II supernova (2006gq), most in the magnitude range g = 21–22, have been reported by J. Frieman, on behalf of the Sloan Digital Sky Survey II collaboration, with details published on CBETs 637 and 643. Also, five additional new apparent supernovae (tabulated below) have been reported from unfiltered CCD images: 2006gr by R. R. Prasad and W. Li (LOSS/KAIT; cf. IAUC 8750); 2006gs by K. Itagaki (cf. IAUC 8751; communicated by S. Nakano, Sumoto, Japan); 2006gt and 2006gy by R. Quimby and P. Mondol (cf. IAUC 8736; details on CBETs 641 and 644); and 2006gz independently by T. Puckett and A. Pelloni (cf. IAUC 8751; source of tabulated data below) and by D. Winslow and Li (LOSS/KAIT).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006gr Sept. 21.20 22^{h}32^{m}22.^{s}67 +30^{o}49′43′′.3 18.8 23′′.2 W, 24′′.5 S
2006gs Sept. 22.80 11 56 08.50 +55 23 29.8 17.0 11′′ E, 3′′ N
2006gt Sept. 18.26 0 56 17.30 − 1 37 46.0 18.5 10′′.8 W, 13′′.5 S
2006gy Sept. 18.30 3 17 27.06 +41 24 19.5 15.0 2′′.0 W, 0′′.4 N
2006gz Sept. 26.01 18 10 26.33 +30 59 44.4 17.6 12′′.5 W, 27′′.6 S

Additional KAIT magnitudes for 2006gr in UGC 12071: Sept. 3.23 UT, [19.5; 22.23, 18.3. SN 2006gr is a type-Ia supernova, well before maximum light on Sept. 23.2 (details on CBET 642). Additional magnitudes from Itagaki for 2006gs in NGC 3977: 2005 May 15, [19.0; 2006 Sept. 21, [17.5; 23.40, 17.0. Nothing is present at the location of 2006gs on Digitized Sky Survey images. Additional approximate magnitudes for 2006gt: Sept. 22.25, 18.2; 23.25, 17.8. SN 2006gt was found by subtracting a co-addition of images taken between 2005 July 16 and Nov. 7 (limiting mag ∼ 19.1) from the 2006 September images. SN 2006gt is a type-Ia supernova, around maximum light on Sept. 23–25 (details on CBETs 641, 645). Additional approximate magnitudes for 2006gy in NGC 1260: Sept. 23.28, 14.8; 24.29, 14.9. SN 2006gy was found by subtracting a co-addition of images taken between 2005 Sept. 28 and Oct. 25 (limiting mag ∼ 18.1) from the 2006 September images; 2006gy may be a type-II(n) supernova or an outburst of an active galactic nucleus (cf. CBETs 647, 648). Additional approximate magnitudes for 2006gz in IC 1277: Sept. 16.21, [19.0 (KAIT); 19, [19.4 (Puckett); 21.12, [18.1 (KAIT); 26.12, 18.0 (KAIT); 27.03, 17.5 (Puckett). Winslow and Li provide the following LOSS position end figures for 2006gz: 26^{s}.33, 43′′.0.

2006 September 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green